Tall Trees, Inc. is using the modified internal rate of return (MIRR) when evaluating projects. The company is able to reinvest cash flows received from the project at an annual rate of 11.30 percent. What is the MIRR of a project if the initial costs are $1,397,200 and the project life is estimated as 9 years? The project will produce the same after-tax cash inflows of 594,400 per year at the end of the year.
Answer #1

Future value = Annuity * [(1 + r)n - 1] / r

Future value = 594,400 * [(1 + 0.113)9 - 1] / 0.113

Future value = 594,400 * [2.620936 - 1] / 0.113

Future value = 594,400 * 14.34457

Future value = $8,526,412.247

MIRR = (Future value / initial investment)1/n - 1

MIRR = (8,526,412.247 / 1,397,200 )1/9 - 1

MIRR = (6.102499)1/9 - 1

MIRR = 1.2226 - 1

MIRR = 0.2226 or 22.26%